- Project Firefly
- Project Vibrio
- Gibson Assembly
- Future applications
- Our Partners
Week 7: Monday 23rd - Sunday 29th August
Today the whole team was in the lab at the same time - it was very exciting.
Now that we have all our oligos, Anja and Will have been assembling the lux operon with a pBAD promoter from V. fischeri using Gibson. They set a race going between the 3 different PCR machines but unfortunately they all failed for different reasons.
Paul has been working on the firefly luciferase from the registry, performing a colony PCR in order to isolate the promoter and luciferase from previous experiments to make the biobrick for submission.
Peter, Ben and Emily (who was now back from Wales) worked on the plate reader experiment that Ben and Peter had planned last week. They plated out some Top10 glycerol stocks of the promoter, rbs and registry luciferase to make liquid culture tomorrow and start the plate reader experiments on Wednesday.
Hannah and Theo also transformed the pJS555 plasmid from Jim Slock, which doesn't contain the phage DNA present in pHK555. Bill continued with his amazing python program for designing primers.
While waiting for the PCR machines to finish the second time round, we practised typing with our faces (this was typed with my nose).
Today Anja and Will continued working with the bacterial luciferase. The transformation had shown that the gibson had not worked, so we needed to figure out what had gone wrong.
After Theo spending a long time with his camera in the dark room, we have got some very cool photos of everyone on the team with bright phosphoreum plates. We also now have a lot of team photos while everyone was here because this might be the last time now until the jamboree - quite a scary prospect. We have some very cool sequence photos of people jumping off steps outside the Sedgwick museum.
Will and Anja also discovered a possible reason that their Gibson didn't work. One of the reverse primers can anneal in the wrong place with its secondary hairpin loop structure, so it seems it had annealed in the wrong place. However, they discovered that the Gibson Master Mix does work which is a relief.
Hannah and Emily continued trying to plan the experiments for when the DNA2.0 order arrives, but there is a lot to think about. Ben and Peter also carried on sorting out the plate reader experiments and should be ready to start the first one tomorrow.
Bill and Anja did something ...
Today we had another lab meeting (finally) at 12pm. Gos, PJ and Shuna were encouraging, which was nice. They were also impressed by Bill's Gibthon program, which he continued to work on as well as helping Ben and Peter in the afternoon. Will and Anja carried on trying to make Gibson work, including designing more primers and running some more gels, while Ben and Peter made the final preparations before their plate reader experiment could start.
In the afternoon the plate reader experiment started at about 3.30pm and should run for 6 hours (all being well). Emily carried on updating the online lab book and notebook and started the experiment for transferring registry luciferase with rbs and promoter into pSB1C3 but the plasmid purification didn't work properly so will do it again tomorrow.
In other news, our order of new E-gels arrived today, which was very unexpected. So we can gel away to our hearts content!
Week 7: Monday 23rd - Sunday 29th August
51. Experiment: Transformation with pJS555 & pSB1C3 (Theo and Hannah)
Received a new plasmid from Jim Slock, so have transformed some chemically competent cells with it to build up a stock and test it later/we think it will glow for longer than the previous one (says Mr. Lovely)).
Theo transformed TOP10 with pSB1C3 from registry.
52. Experiment: Check for promoter+rbs+p.p.luc - obtain from the colony from transformation on page 37. (Paul)
Take sample from colony with loop and put in 20μl of water to obtain DNA template.
Then prepare on isoFreeze PCR chiller in lister order:
1. 98°C for 15 mins
2. 98°C for 10 secs
3. 65°C for 30 secs
4. 72°C for 1 min
5. repeat 2-4 35 times
6. 72°C for 10 mins
7. 4°C forever
Tube 1: 197.4ng/μl
Tube 2: 235.3ng/μl
53. Experiment: Extracting glycerol stocks of TOP10 with promoter+rbs+p.p.luc from registry. (Emily and Peter)
1. Take tube of glycerol stock from -80°C 2. Use loop/tooth pick (we used a loop) to streak on ampicillin resistance plate. Let it melt first, then streak.
NB NEVER let glycerol stocks thaw. They should be out and back in the -80°C freezer in 2-3mins max
3. Place in 30°C incubator overnight.
To be continued tomorrow.
54. Check for promoter+rbs+luc from p32 by luciferin injection. (Paul and Ben)
We used protocol described on p22.
3x Freezymes at -80°C, thawing at 37°C
55. Experiment: Gibson assembly of pBAD luxCDABEG in psB1C3
Added (on ice!) in 5 individual (A-E) PCR tubes
Programs on PCR machines
Results: Machine 1 had a power fail very close to the end of the program Machine 2 had a loose lid; the tubes popped open and there was no liquid left Machine 3 had entirely melted the top of the PCR tube
The experiment was repeated, but this time all reactions (A-E) were run in Machine 2 (see p.38 for program).
blanked with distilled water
Gel loading mixtures were made up according to:
final volume 20µl
A 1% agarose E-gel was loaded as follows
Bands were observed at the following lengths:
Bands for pA, B, pC, D & pE were cut out the gel and DNA was extracted following the QIAquick Gel Extraction Kit protocol
Preparation of Gibson assembly 1.33x master mix
Gibson Assembly Reaction added to PCR tube:
Incubated for 1h at 50C Stored at 4C overnight
56. Experiment: Transformation of TOP10, red strain and with Gibson assembly (pBAD, luxCDABEG, pSB1C3) (Will and Anja)
Followed protocol on p13+14. Transformed:
Plated 150µl on LB agar plates with Chl. each (9pl.)
57. Experiment: Repeat gel electrophoresis from p.39+40 (Will and Anja)
Followed protocol on p.39+40 (pcr products had been kept in -20°C overnight).
Results: A, C, D, pA & pC were of appropriate size. B (again 3-5kb) and E, pE (slightly lower than expected, closer to D than C) were of inappropriate size.
It was decided to run new PCR reactions for B (CD) and E (pSB1C3) following the protocol on p.38.
Nanodrop measurements: (blanked with Di H20)
B (CD) 980.7ng/µl
E (pSB1C3) 127.5ng/µl
Gel loading mixtures made up as follows:
A 1% Agarose (E-gel) was loaded as follows:
58. Experiment: Making liquid culture of tetR repressed promoter, rbs + luc in TOP10 (from registry) (Peter and Emily)
Make LB broth with campicillin, adding 200µl ampicillin to 200ml LB to make 100µg/ml concentration final solution.
Plates prepared from glycerol stocks had grown well, but were transferred from 30°C to 37°C incubator for better growth - the temperature effect on luciferase won't matter. The coonies were very small. A single colony was chosen using a loop, placed in 3ml of LB&Amp and shaken. 3 tubes were used and placed in the 30°C incubator overnight.
59. Experiment: Gel Extraction (Anja)
Results from Gel electrophoresis on p.42.
E showed a band slightly above 2kb.
B again showed a band between 3-5kb.
Bands for A & pA, B, pC, D, pE were cut out from 'repeated' gel on p.41 and 'Enew' as well as a trace of seemingly correctly sized B ('B trace') were cut out from gel on p.42. DNA was extracted following the 'QIAquick Gel Extraction Kit (miniElute columns)' protocol.
60. Experiment: PCR of B trace with primers for amplification of B & gel electrophoresis (Will and Anja)
Followed protocol on p.38.
Nanodrop measurements: B trace PCR 492.2ng/µl
Gel loading mix:
Gel loading as follows:
61. Experiment: Diagnostic gel of Gibson assembly reaction (Bill and Anja)
Gel loading mixtures:
Result: In neither G1, G2 or G3 did any high molecular weight bands occur other than those for pA, B, pC, D and pE individually => Gibson assembly reaction did not work.
62. Experiment: Test Gibson Master Mix for functionality (Bill and Anja)
Gibson assembly reaction
Added to a PCR tube:
Incubated for 1h at 50°C
(All other wells were filled with DI H20)
Gel loaded as follows:
=> Gibson Master Mix is functional
63. Experiment: Gibson assembly of pBAD, luxCD AB EG and pSB1C3 as well as all individual 'neighbour parts' in pairs (Will and Anja)
j1-j6: A->E (pA, B, pC, D & pE) refer to PCR reactions (that have not been run on a gel and gel extracted)
2 separate gels were run:
1. Gel loading mixtures:
Gel was loaded as follows:
2. Gel loading mixtures:
Gel was loaded as follows:
Transformation of TOP10 with 1)G1 and 2)j1
Followed protocol on p13+14, however, we transformed with 10µl Gibson reaction instead of 2µl. Plated 150µl on LB agar plates with Chl. Incubated at 30°C overnight.
Result: All 4 plates blank after overnight :( All 4 growth after 2 days, RFP present, no glow.
Results: No colonies observed from above transformation.
64. Expt: Transfer of promoter+rbs+P.P.Luc into pSB1C3 for submission to registry (Emily + Anja)
Followed protocol on p31 (nanodrop:[plasmid]=26.5ng/µl)
Restriction enzyme digest
Prepated at RT in the following order:
The digest was prepared twice. It was checked that 15µl of plasmid DNA would not contain more than 1µg of DNA. The reaction was mixed gently, spun down and incubated at 37°C for 4h45min (in PCR machine).
Nanodrop measurements: Digest 1 -> 18.3ng/µl Digest 2 -> 22.8ng/µl
pSB1C3 supplied at 25ng/µl. 10-100ng of linearised vector DNA should be added to ligation mix. 3µl give 75ng. 3:1 molar excess of insert DNA should be added.
75ng/2035b x 1727b x 3 = 190ng => 9µl
To a PCR tube add:
The ligation was prepared twice. Mix was vortexed, spun down and incubated at 22°C for 1h45min (in PCR machine).
Followed protocol on p13+14. Transformed 2x50µl TOP10cc with ligation 1 and ligation 2 (10µl of ligation rather than 2µl). Plated on LB agar plates with Chloramphenicol (plated 150µl).
65. Expt: Gibson assembly of pBAD, luxCDABEG (amplified in one fragment) and pSB1C3 (Will and Anja)
(BBa_J04450 had one TOP10 colony in DI H20)
PCR mix made up as on p38.
Program of PCR machine:
For pBAD and luxCDABEG:
1% (self-cast) Agarose gel loaded as follows:
+ 5µl 6x orange LD (mixed before adding to gel) to everything except the ladder.
run at 120V
Results: Only for pBAD did we observe a band (this was of correct size).
=> PCR of luxCDABEG from pJS555 failed
=> PCR of pSB1C3 from BBa_J04450 transformed colony failed.
Following the 'MinElute Gel Extraction' protocol pBAD was isolated (using MinElute columns) and stored at 4°C.
66. Expt: PCR of pJS555 colony PCR of cells containing pSB1C3 (&RFP biobrick), PCR of pBAD (Will & Hannah)
1. pJS555 was diluted 10x in nuclease free water (taken from vial sent by James Slock), and PCRed in 3 fragments: CD, AB, EG using primers ordered on 17.08.10.
2. A colony grown up with a pSB1C3 plasmid (with RFP biobrick) was suspended in 20µl DI autoclaved water. 2µl of this was added to standard PCR mix as "template".
3. Purified I0500 (from Shuna)
Naming system is as on pp38-41.
A, C, b, d in machine (2)
e in machine (3)
In addition, another b was run in machine (3) in an attempt to get rid of the problematic secondary structure in the template DNA (hypothesised). This tube was labelled B72.
After PCR tubes were run on Agarose on a single gel
Machine (2) did not have top screwed on properly so tube containing D was lost. For the purposes of this assembly a aprevious PCR product of D was used, though there was only 10µl of this product left, so a reduced amount was used.
The layour was: EZ ladder II, A, B, C, D, E, B72
After running for an hour, no banding was produced by E and B72.
But good bands of A, B, C, D were produced, including the CORRECT SIZED B BAND! :) :)
These 4 bands were extracted and purified according to Qiagen "Qiagen gel extraction kit", final ste 10µl of EB was added, although a normal column was used.
67. Expt: Gibson assembly of above fragments (Will)
A previously gel extracted E was used to replate the failed band (p41 original E).
2µl A, B, C, D, E
+30µl 1.33x Gibson master mix was added.
Incubated at 50°C for 60m.
68. Expt: Transformation of TOP10 with Gibson product above (Will)
Followed protocol on pp13-14.